KTeV OWL Shift - 8/11/99 (Wed.)

Earl C. Swallow (earls@elmhurst.edu)
Wed, 11 Aug 1999 07:01:47 -0500 (CDT)

Shift Summary for OWL shift, 08/11/99 (Wed.)

On Shift: Sydney Taegar & Earl Swallow + BobT & RickK

Hours of beam available at all intensities: 0.0 (out of 7.0)
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

Accelerator down period continues.


========================================================================
Runs ended this shift:
14179 - cosmic ray
14180 - laser scan
14181 - cosmic rays
14182 - pedestals
14183 - cosmic rays
14184-87 - scrap

Tapes brought to FCC this shift: NONE

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

VERY Quiet shift with no beam -- UGH!

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:00 Hall is secure. BobT & RickK are on access.
They return shortly after start of shift, having
effected a successful repair.
Page #19 in KTeV Log XXI.

00:41 BobT and RickK leave us with all of the excitement.

00:45 CsI #1691 gain updated to correct high (28%)
deviation in gain drift plot. Start cosmic Run
14179 - #1691 looks good now.

01:03 Start laser scan - Run 14180.

01:40 Only one copy of second sheet of Gas Check List
available -- I fake it for now and send email
requesting more copies.

01:48 End Run 14180 - Good laser scan. Start cosmic
Run 14181,

03:15 End Run 14181.

03:26 Start pedestal Run 14182.

03:30 Cycle analysis magnet (NM4AN) to get low residual
field and POS polarity (reversed from previous NEG
running) for MUON runs when beam returns. (This
gives -0.5 Gauss rather than -1.7 Gauss from simple
run-down.) After MUON running, just run it up with
AN832UP script.

03:32 End pedestal Run 14182 and updated pedestals.

03:40 Start cosmic ray Run 14183.

04:20 Repeatedly losing permits in stream 5. Multiple
KFRENDS and power cycling laser crate fixes it.

04:40 End cosmic ray Run 14183.

05:30 ST tested new version of DOT which checks tape
labels before appending. Seems to work fine, but
will not be installed yet for the sake of safety.

06:00 Start cosmic ray Run 14188 with OLD DOT version.

06:30 ST has been recopying incorrectly labeled tapes.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Try to get us some beam.
---> Load QK*249-251 and QKB087 and APPEND to these tapes.

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Professor & Chairperson                    (or)   earls@fnal.gov
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)